# An Introduction to Zero-Order Optimization Techniques for Robotics

### v1

@@ I Introduction
In recent years, zero-order (or derivative-free) optimization techniques [1] have gained a lot of popularity in the robotics community. While zero-order optimization is a well-established field, its widespread deployment in robotics has only been made possible by the recent advances in parallel computing and GPU hardware. These improvements have enabled the use of zero-order Trajectory Optimization (TO) techniques in a Model Predictive Control (MPC) fashion on complex robots [2, 3]. In parallel, Reinforcement Learning (RL) has emerged as a powerful tool and demonstrated state-of-the-art capabilities in tasks such as locomotion or manipulation [4, 5].
@@
While these approaches may appear unrelated at first, they share a common property: they do not require access to the simulator’s gradients. This characteristic allows them to optimize non-smooth objective functions, which typically arise in problems involving contact such as locomotion or manipulation. Another benefit is that it avoids the tedious development of efficient gradient implementations. Furthermore, while deterministic gradient-based algorithms, such as gradient descent or Newton method, are prone to getting stuck in local minima, zero-order techniques are mostly stochastic, which can help escape local minima.
@@
Zero-order optimization has a long history [6]. In its most generic form, zero-order optimization performs random search by generating samples at random and updating an estimate of the optimal solution based on the performance of each sample. In the 1970s, Evolutionary Strategies (ES) gained popularity [7] and later culminated with the CMA-ES algorithm [8]. In TO, predictive sampling [9] is a well-known and simple random search technique. Another popular algorithm is MPPI [10], which was initially derived from information theory. In RL, Reinforce [11] was one of the first major algorithms proposed for continuous action space problems. These algorithms have since inspired many variations.
@@
While most of these algorithms have strong theoretical foundations, this is not necessarily well known in the robotics community. One reason for this is that these algorithms often fall at the intersection of many fields such as optimization, statistics, machine learning and control. By looking at results from these different fields, we aim to answer two questions: 1) What are the convergence guarantees of the algorithms used in robotics? 2) What are the theoretical tools explaining the ability of these algorithms to escape local minima?
@@
Recently, [12] introduced Gaussian smoothing and understood random search from an optimization perspective. Inspired by this work, we propose a unified perspective to understand the zero-order optimization techniques popular in the robotics community. More specifically, we are interested in algorithms that can find solutions to the following problem:
$\min_{x\in\mathbb{R}^{n}}f(x)$ (1)
@@
This formulation encompasses a wide range of problems encountered in robotics, such as TO and RL. Both of these techniques define a high-level goal through a cost function that has to be minimized (or a reward function that has to be maximized). However, they differ in the optimization space. TO optimizes directly over trajectories, whereas RL optimizes over policies. Furthermore, they differ in the way they are deployed. These algorithms can be used either online (i.e. executed at runtime) or offline. On one hand, TO can either be solved online in an MPC loop or solved offline to generate data to train a policy. On the other hand, RL is mostly solved offline. While these differences may seem fundamental, we show that the algorithms used to solve these problems share many conceptual similarities.
Related work
@@
[13] first showed the similarities between MPPI and CMA-ES. More recently, [14] understood MPPI as an approximate gradient method and [3] connected MPPI to diffusion models. Our work builds on these insights and further investigates explanations for the success of MPPI. In particular, we shed light on the connection between MPPI and recent results showcasing the benefits of the log-sum-exp transform [15]. Lastly, we benchmark several state-of-the-art approaches on various robotic examples.
@@
[16, 17] first made the connection between RL and random search. These works used this insight to derive TO algorithms based on random search. In contrast, we study how to leverage this understanding to derive novel RL algorithms.
Contribution
@@
In this work, we propose a mathematical introduction to zero-order optimization algorithms used in robotics. This unified perspective provides a simple way to understand techniques such as MPPI, Covariance Matrix Adaptation (CMA) or RL policy gradient methods. In addition, we show how this unified view allows us to naturally derive novel competitive methods. Lastly, we discuss the theoretical concepts explaining why stochastic algorithms are well-suited to avoid getting stuck in local minima.
Outline
@@
Section II introduces the main concepts of random search. Section III connects popular TO algorithms to random search. Section IV investigates the connections between Gaussian smoothing and RL. Lastly, Section V discusses how to improve the sample efficiency of these algorithms by relying on a population of samples.
@@ II-A Simple random search
In this first section, we present how to use a population of samples (or particles) to search for a global solution. We divide the population search algorithms in two broad classes: global search and local search algorithms.
II-A1 Global Search
Global search can be considered as the simplest global optimization technique. The idea is to iteratively sample a random point. If the sample has a value lower than the previous one, it is kept as the current best guess. Otherwise, it is disregarded. Algorithm 1 summarizes the procedure.
@@ 5             $x\leftarrow\tilde{x}$
Algorithm 1 Simple Global Search
Interestingly, the sampling distribution of Algorithm 1 ignores all the previous estimates. For this reason, it can be referred to as a global algorithm. Under some mild conditions on the objective function, Algorithm 1 converges to a global minimum [18]. Unfortunately, in practice, this method performs poorly as it is subject to the curse of dimensionality. That is to say that the number of function evaluations required to find a global solution grows exponentially with the dimension of the problem.
@@ II-A2 Greedy Local Search
A common attempt to break the curse of dimensionality is to exploit problem-specific knowledge. For instance, exploiting previous information about the shape of the function can help design an empirically more efficient algorithm. This is called Greedy Local Search. The idea is to iteratively search around the previous estimate, with additive Gaussian noise, for instance. Then, as in the previous algorithm, the sample is kept only if it improves the function value. Algorithm 2 summarizes the procedure.
Input: $x_{0}\in\mathbb{R}^{n}$
@@ 1 $x\leftarrow x_{0}$
       Sample $d$ in $\mathbb{R}^{n}$ // with a centered Gaussian distribution
@@ 4             $x\leftarrow x+d$
By design, Algorithm 2 depends on previous estimates. For that reason, it can be referred to as a local algorithm. However, this notion should not be confused with the concepts of local and global solutions. In fact, although the convergence guarantees of Algorithm 2 are more restrictive than those of Algorithm 1, it is still possible to prove convergence of Algorithm 2 to a global solution under some regularity conditions [18]. Consequently, both local and global search can have global convergence guarantees. However, while both are subject to the curse of dimensionality, local search techniques can be more effective empirically. In fact, as we will see in the next section, it is well-known that local search techniques relying on gradient information can exhibit strong convergence rates and guarantees in convex settings.
@@ II-B Random search via gradient approximation
In this section, we present a few ways to perform approximate gradient descent using only function evaluations. The class of methods under study is summarized in Algorithm 3.
@@ 3       $x\leftarrow x-\alpha g$
Here, $\alpha$ is the step size (or learning rate). There exist many ways to estimate the gradient with multiple function evaluations [19]. Among the most widely used ones is the forward finite differences technique:
$\sum_{j=1}^{n}\frac{f(x+\mu e_{j})-f(x)}{\mu}e_{j}$ (2)
where $e_{j}\in\mathbb{R}^{n}$ represents the $j^{th}$ canonical vector. Alternatively, one may use central finite differences. While this estimate can be very effective, it requires $n+1$ function evaluations, which is a limitation in high-dimensional settings.
@@ II-B1 Random Coordinate Descent
$g=\frac{f(x+\mu e_{j})-f(x)}{\mu}e_{j}$ (3)
where $j$ can be chosen uniformly at random in $\{1,\dots,n\}$. Note that $g$ depends on $x$.
Remark 1.
In the convex case and without finite difference approximation, Random Coordinate Descent requires at most $n$ times more steps compared to classical gradient descent [20]. In addition, on a certain class of functions, it can require as many steps as gradient descent while requiring $n$ times less function evaluations.
Random Coordinate Descent samples discrete coordinates, but it is also possible to sample along arbitrary directions.
@@ II-B2 Simultaneous Perturbation Stochastic Approximation (SPSA)
$g=\frac{f(x+\mu\Delta)-f(x-\mu\Delta)}{2\mu}\Delta$ (4)
where each component of $\Delta$ is equal to ±$\pm 1$ according to independent Bernoulli distributions. Note that Spall derived SPSA using central differentiation instead of the forward finite difference. Under reasonable assumptions, SPSA can reach similar performances as finite differences while requiring $n$ times less samples [18].
@@ II-B3 Gaussian Smoothing
In [12], Nesterov proposes another approach that samples the gradient of a smoothed surrogate function
$f_{\mu}(x)=\mathbb{E}[f(x+\mu u)]$ (5)
where $u\sim\mathcal{N}(0,\Sigma)$. This is Gaussian smoothing, commonly referred to as randomized smoothing. As $\mu$ goes to zero, $f_{\mu}$ tends to $f$. As a result, for small values of $\mu$, the gradients of the surrogate function $f_{\mu}$ are close to those of the original function $f$. The advantage of this surrogate is that its gradient can be evaluated via an average of function evaluations. To understand this, let’s first define $\kappa=\sqrt{(2\pi)^{n}\det(\Sigma)}$, the normalization factor of the multivariate Gaussian distribution. With a change of variables, we can write:
[…]
Consequently,
[…]
$=\mathbb{E}\left[\frac{1}{\mu}f(x+\mu u)\Sigma^{-1}u\right]$
$=\mathbb{E}\left[\frac{f(x+\mu u)-f(x)}{\mu}\Sigma^{-1}u\right]$ (7)
The last equality comes from the fact that $\mathbb{E}\left[f(x)\Sigma^{-1}u\right]=0$. Hence, the gradient of $f_{\mu}$ can be estimated using only function evaluations of the original function, $f$ (at the cost of approximating an expectation with samples). Note that the term $\Sigma^{-1}u$ is the gradient of the logarithm of the probability density function with respect to the mean; this is often called the log-likelihood trick. This is sometimes written:
$\nabla f_{\mu}(x)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\nabla_{x}\log(p_{x}(z))\right]$ (8)
where $p_{x}$ is the density of the Gaussian distribution with mean, $x$, and with covariance, $\Sigma$. This alternative view will be useful to understand CMA. In the general case, the expectation is intractable, however, one can rely on sampling to estimate it. For instance, one can use the following estimate in Algorithm 3:
$g=\frac{f(x+\mu u)-f(x)}{\mu}\Sigma^{-1}u$ (9)
While $\frac{1}{\mu}f(x+\mu u)\Sigma^{-1}u$ is a valid gradient estimate, it can be arbitrarily large. In contrast, the estimate (9) is invariant to constant translations of the function. Intuitively, this will help reduce the variance of the estimate. Note that it is also possible to use an estimate with central differentiation:
$g=\frac{f(x+\mu u)-f(x-\mu u)}{2\mu}\Sigma^{-1}u$ (10)

### v2

@@ I Introduction
In recent years, zero-order (or derivative-free) optimization techniques [1] have gained a lot of popularity in the robotics community. While zero-order optimization is a well-established field, its widespread deployment in robotics has only been made possible by recent advances in parallel computing and GPU hardware. These improvements have made it possible to deploy sampling-based Model Predictive Control (MPC) on complex robotic systems [2, 3]. In parallel, Reinforcement Learning (RL) has emerged as a powerful tool and has demonstrated state-of-the-art capabilities in locomotion or manipulation [4, 5].
@@
While these approaches may appear unrelated at first, they share a common property: they do not require access to the simulator’s gradients. This characteristic allows them to optimize non-smooth objective functions, which typically arise in problems involving contact, such as locomotion or manipulation. Another benefit is that it avoids the tedious development of efficient gradient implementations. Furthermore, while deterministic gradient-based algorithms, such as gradient descent or Newton’s method, are prone to getting stuck in local minima, zero-order techniques are mostly stochastic, which can help escape local minima.
@@
Zero-order optimization has a long history [6]. In its most generic form, zero-order optimization performs a random search by generating samples at random and updating an estimate of the optimal solution based on the performance of each sample. In the 1970s, Evolutionary Strategies (ES) gained popularity [7] and later culminated with the CMA-ES algorithm [8]. In TO, predictive sampling [9] is a well-known and simple random search technique. Another popular algorithm is MPPI [10], which was initially derived using information-theoretic arguments. In RL, Reinforce [11] was one of the first major algorithms proposed for continuous action space problems. These algorithms have since inspired many variations.
@@
While most of these algorithms have strong theoretical foundations, this is not always widely recognized within the robotics community. One reason for this is that these algorithms often lie at the intersection of many fields, such as optimization, statistics, machine learning, and control. By bringing together results from these areas, our goal is both to bring to light the theoretical properties of existing algorithms and to leverage this understanding to design novel algorithms for robotics.
@@
Recently, [12] introduced Gaussian smoothing and understood random search from an optimization perspective. Inspired by this work, we propose a unified perspective to understand the zero-order optimization techniques that are popular in the robotics community. More specifically, we are interested in algorithms that can find solutions to the following problem:
$\min_{x\in\mathbb{R}^{n}}f(x).$ (1)
@@
This formulation encompasses a wide range of problems encountered in robotics, such as TO and RL. Both of these techniques define a high-level goal through a cost function that has to be minimized (or a reward function that has to be maximized). However, they differ in the optimization space. TO optimizes directly over trajectories (i.e. a finite-dimensional sequence of vectors), whereas RL optimizes over policies (i.e. functions, which are infinite-dimensional objects). Furthermore, they differ in the way they are deployed. These algorithms can be used either online (i.e. executed at runtime) or offline. On one hand, TO can either be solved online in an MPC loop or solved offline to generate data to train a policy. On the other hand, recent RL successes relied heavily on offline training in simulation [4, 5]. While these differences may seem fundamental, we show that the algorithms used to solve these problems share many conceptual similarities.
@@
In the trajectory optimization community, several works have studied derivative-free algorithms and established connections among them. For instance, [13] first showed the similarities between MPPI and CMA-ES. Later, [14] understood MPPI as an approximate gradient method and [3] connected MPPI to diffusion models. Our work builds on these insights and proposes a unifying framework to understand these algorithms. In addition, we shed light on the connection between MPPI and recent results showcasing the benefits of the log-sum-exp transform [15]. Lastly, we benchmark several state-of-the-art approaches on various robotic examples.
@@
In the RL community, [16, 17] first explored the use of random search in the parameter space of policies as an alternative to traditional RL algorithms. In this work, we instead investigate how random search techniques can explain the success of popular RL algorithms. More recently, [18, 19] drew the connection between policy gradient techniques and Nesterov’s random search [12]. However, these works used this insight to derive TO algorithms based on random search. In contrast, we study how to leverage this understanding to derive novel RL algorithms. Lastly, we refer the reader to [20, 21] for comprehensive surveys on the connection between RL and ES techniques.
@@
Overall, a broad perspective connecting gradient-free approaches in TO and RL is missing. To bridge this gap, we propose a mathematical introduction to zero-order optimization algorithms used in robotics. This unified treatment provides a simple way to understand techniques such as MPPI, Covariance Matrix Adaptation (CMA), and RL policy gradient methods. In addition, we show how this unified view allows us to naturally derive novel competitive methods as a byproduct. Lastly, we discuss the theoretical concepts explaining why stochastic algorithms are well-suited to avoid getting stuck in local minima.
@@
Section II introduces the main concepts of random search and sampling-based gradient approximation. Section III connects popular TO algorithms to random search. Section IV investigates the connections between Gaussian smoothing and RL. Section V discusses how to improve the sample efficiency of these algorithms by relying on a population of samples. Lastly, Section VI briefly reviews existing parallel computing libraries.
@@ II Random search
In this section, we introduce how a population of samples (or particles) can be used to search for global solutions. The algorithms presented here will serve as building blocks for the understanding of widely used TO and RL algorithms. Depending on the context, the samples may correspond to trajectories in TO or to policy parameters in RL.
@@ II-A Simple random search
First, we study two basic random search approaches: global search and local search algorithms. Although rarely used in practice, their convergence guarantees help to understand more sophisticated algorithms.
II-A1 Pure (or Global) Random Search
The most naive random search approach consists of iteratively sampling points at random. If the sampled point has a value lower than the current best, it is kept as the new best guess; otherwise, it is discarded. This method, known as Pure Random Search, is summarized in Algorithm 1.
@@ 5    $x\leftarrow\tilde{x}$
Algorithm 1 Pure Random Search
Interestingly, the sampling distribution of Algorithm 1 ignores all the previous estimates. For this reason, it is sometimes referred to as a global algorithm. Under some mild conditions on the objective function, Algorithm 1 converges to a global minimum [22]. Unfortunately, in practice, this method performs poorly as it is subject to the curse of dimensionality. That is to say, the number of function evaluations required to find a global solution grows exponentially with the dimension of the problem.
@@ II-A2 Greedy Local Search
A common attempt to break the curse of dimensionality is to exploit problem-specific knowledge. For instance, exploiting previous information about the shape of the function can help design an empirically more efficient algorithm. This is called Greedy Local Search. The idea is to iteratively search around the previous estimate with additive Gaussian noise, for instance. Then, as in the previous algorithm, the sample is kept only if it improves the function value. Algorithm 2 summarizes the procedure.
Input: $x_{0}\in\mathbb{R}^{n}$, $\Sigma$
@@ 1 $x\leftarrow x_{0}$
3   Sample $d$ with $\mathcal{N}(0,\Sigma)$
@@ 5    $x\leftarrow x+d$
By design, Algorithm 2 depends on previous estimates. For that reason, it can be referred to as a local algorithm. However, this notion should not be confused with the concepts of local and global solutions. In fact, although the convergence guarantees of Algorithm 2 are more restrictive than those of Algorithm 1, it is still possible to prove the convergence of Algorithm 2 to a global solution under some regularity conditions [22]. Consequently, both local and global searches can have global convergence guarantees. However, while both are subject to the curse of dimensionality, local search techniques can be more effective empirically. In fact, as we show next, it is well known that local search techniques relying on gradient information can exhibit strong convergence rates and guarantees in convex settings.
@@ II-B Random search via gradient approximation
In this section, we present various ways to perform approximate gradient descent using only function evaluations. In later sections, we show how widely used algorithms in TO and RL can be derived from the basic principles introduced here. The class of methods under study is summarized in Algorithm 3.
@@ 3 $x\leftarrow x-\alpha g$
Here, $\alpha$ is the step size (or learning rate). There are many ways to estimate the gradient with multiple function evaluations [23]. Among the most widely used techniques is the forward finite difference technique:
$\sum_{j=1}^{n}\frac{f(x+\mu e_{j})-f(x)}{\mu}e_{j},$ (2)
where $e_{j}\in\mathbb{R}^{n}$ represents the $j^{th}$ canonical vector and where $\mu>0$ is a scalar. Alternatively, one may use central finite differences. While this estimate can be very effective, it requires $n+1$ function evaluations, which is a limitation in high-dimensional settings.
@@ II-B1 Random Coordinate Descent
$g=\frac{f(x+\mu e_{j})-f(x)}{\mu}e_{j},$ (3)
where $j$ can be chosen uniformly at random in $\{1,\dots,n\}$. Note that $g$ depends on $x$. In the convex setting and without finite-difference approximation (i.e. with the exact gradient), Random Coordinate Descent requires, on average, at most $n$ times more steps compared to classical gradient descent [24]. In addition, for a certain class of functions, it can require as many steps as gradient descent while requiring $n$ times fewer function evaluations. Thus, this stochastic estimate offers the hope of matching the gradient’s descent performance with a few samples. Random Coordinate Descent samples discrete coordinates, but it is also possible to sample along arbitrary directions.
@@ II-B2 Simultaneous Perturbation Stochastic Approximation (SPSA)
$g=\frac{f(x+\mu\Delta)-f(x-\mu\Delta)}{2\mu}\Delta,$ (4)
where each component of $\Delta$ is equal to $\pm 1$ according to independent Bernoulli distributions. Note that Spall derived SPSA using central differentiation instead of the forward finite difference. Under reasonable assumptions, SPSA can achieve performance similar to that of finite differences while requiring $n$ times fewer samples [22]. Hence, similarly to Random Coordinate Descent, this stochastic estimate offers the hope of matching the gradient’s descent performance with a few samples.
While both approaches lead to effective algorithms, the next technique provides a more insightful interpretation.
@@ II-B3 Gaussian Smoothing
[12] propose another perspective. Instead of directly approximating the gradients of the original function, they propose studying a smoothed surrogate function
$f_{\mu}(x)=\mathbb{E}[f(x+\mu\epsilon)],$ (5)
where $\epsilon\sim\mathcal{N}(0,\Sigma)$. This is Gaussian smoothing, commonly referred to as randomized smoothing (RS). As $\mu$ approaches zero, $f_{\mu}$ tends to $f$. As a result, for small values of $\mu$, this surrogate only slightly changes the objective function (and therefore the minimum location) but adds smoothness [12]. Importantly, the advantage of this surrogate is that its gradient can be evaluated via an average of function evaluations. To understand this, let’s first define the constant: $\kappa=\sqrt{(2\pi)^{n}\det(\Sigma)}$, the normalization factor of the multivariate Gaussian distribution. With a change of variables ($z=x+\mu\epsilon$), we can write:
$f_{\mu}(x)$ $=\frac{1}{\kappa}\int_{\epsilon}f(x+\mu\epsilon)e^{-\frac{1}{2}\epsilon^{\top}\Sigma^{-1}\epsilon}\,d\epsilon$
$=\frac{1}{\kappa\mu^{n}}\int_{z}f(z)e^{-\frac{1}{2\mu^{2}}(z-x)^{\top}\Sigma^{-1}(z-x)}\,dz.$ (6)
This allows differentiating $f_{\mu}$ without needing the gradient of the original function, $f$:
$\nabla f_{\mu}(x)$ $=$ (7)
$\frac{1}{\kappa\mu^{n+2}}$ $\int_{z}f(z)e^{-\frac{1}{2\mu^{2}}(z-x)^{\top}\Sigma^{-1}(z-x)}\Sigma^{-1}(z-x)dz.$
With another change of variable ($\epsilon=\mu^{-1}(z-x)$), we have:
$\nabla f_{\mu}(x)$ $=\frac{1}{\kappa\mu}\int_{\epsilon}f(x+\mu\epsilon)e^{-\frac{1}{2}\epsilon^{\top}\Sigma^{-1}\epsilon}\Sigma^{-1}\epsilon\,d\epsilon$
$=\mathbb{E}\left[\frac{1}{\mu}f(x+\mu\epsilon)\Sigma^{-1}\epsilon\right].$ (8)
Hence, the gradient of $f_{\mu}$ can be estimated using only function evaluations of the original function, $f$. However, this comes at the cost of approximating an expectation with samples. Note that the term $\mu^{-1}\Sigma^{-1}\epsilon$ is the gradient of the logarithm of the probability density function with respect to the mean; this is often called the log-likelihood trick. This is sometimes written:
$\nabla f_{\mu}(x)=\mathbb{E}_{z\sim\mathcal{N}(x,\Sigma)}\left[f(z)\nabla_{m}\log(p_{m=x}(z))\right],$ (9)
where $p_{m}$ is the density of the Gaussian distribution with mean $m$ and covariance $\Sigma$. This alternative view is useful for understanding the CMA algorithm, explained in Section III-C. Lastly, since $\mathbb{E}\left[f(x)\Sigma^{-1}\epsilon\right]=0$, we can write:
$\nabla f_{\mu}(x)$ $=\mathbb{E}\left[\frac{f(x+\mu\epsilon)-f(x)}{\mu}\Sigma^{-1}\epsilon\right].$ (10)
In the general case, the expectation is intractable. The idea is then to approximate the gradient of the original function with a stochastic estimate of the gradients of the surrogate. For instance, one can use the following estimate in Algorithm 3:
$g=\frac{f(x+\mu\epsilon)-f(x)}{\mu}\Sigma^{-1}\epsilon.$ (11)
While $\frac{1}{\mu}f(x+\mu\epsilon)\Sigma^{-1}\epsilon$ is a valid gradient estimate, it can be arbitrarily large. In contrast, the estimate (11) is invariant to constant translations of the function. Intuitively, this reduces the variance of the estimate. Note that it is also possible to use an estimate with central differentiation:
$g=\frac{f(x+\mu\epsilon)-f(x-\mu\epsilon)}{2\mu}\Sigma^{-1}\epsilon.$ (12)
Remark 1.
Alternatively, one may sample, $\epsilon$, from the surface of a unit sphere [25]. This technique guarantees that the norm of the direction vectors is bounded.
@@
To summarize, although the assumptions in [22, 24, 12] vary, the underlying ideas are of the same nature; sampling along random directions may match the performance of gradient descent while requiring $n$ times fewer function evaluations.
II-C Simulated Annealing
So far, we have seen that performing random approximate gradient descent can be sample efficient in the convex setting. However, one may ask whether these algorithms can converge to global solutions in nonconvex settings, as it is the case for the Greedy Local Search introduced in Algorithm 2. It is well known that, in its vanilla form, gradient descent is prone to local minima. However, additive Gaussian noise can lead to asymptotic global convergence guarantees. To understand why, we need to study the following Langevin dynamics:
$dX_{t}$ $=\nabla_{x}\log p(X_{t})+dW_{t}.$ (13)
Here $p$ is a probability distribution, and $W_{t}$ is a Brownian motion. $\nabla_{x}\log p$ is the score function of the distribution. Note that the gradient is taken with respect to the input of $p$ (not with respect to the parameters of $p$ as in the log-likelihood trick of Gaussian smoothing). This Stochastic Differential Equation (SDE) has been extensively studied, and it can be shown that $X_{t}$ converges to the distribution defined by $p$ [26]. Consequently, this SDE can be used to sample from a distribution with probability density $p$. One way to leverage this result in the context of global optimization is to choose the following probability distribution:
$p\propto\exp\left(-\frac{1}{\lambda}f\right),$ (14)
$\nabla_{x}\log p(x)=-\frac{1}{\lambda}\nabla_{x}f(x).$ (15)
Then, this SDE can be discretized to derive a recursive algorithm. Following this idea, [28] show that with an appropriate schedule of step sizes $\alpha_{k}$ and $\gamma_{k}$, the discretized Stochastic Langevin Gradient Dynamics,
$x_{k+1}=x_{k}-\alpha_{k}g_{k}+\gamma_{k}\epsilon_{k},$ (16)
guarantees asymptotic convergence to global solutions. Here, $\epsilon_{k}$ is Gaussian noise and $g_{k}$ is of the form $(\nabla f(x_{k})+\xi_{k})$}, with $\xi_{k}$ representing stochastic noise. Hence, $g_{k}$ can be interpreted as a noisy approximation of the gradient. Intuitively, the Gaussian noise $\epsilon_{k}$ helps to escape local minima. The step sizes $\alpha_{k}$ and $\gamma_{k}$ are chosen such that the corresponding SDE has a temperature $\lambda$ that tends slowly to zero. This is called Simulated Annealing. Intuitively, slowly decreasing the temperature guides the iterates towards a global solution.
These results can be used to prove the global convergence of SPSA. More precisely, the iterative scheme in Equation (16) with $g$ as the SPSA estimate defined in Equation (4) converges to a global minimum [29]. In fact, in its default version (i.e. without additive noise), SPSA converges to a global minimum [29]. This is because the noise of the estimate (i.e. $\xi_{k}$) is enough to escape local minima and achieve global convergence. We are not aware of such results in the context of randomized smoothing. However, we can conjecture that similar results may be valid.
To summarize, we have seen that stochasticity can help gradient descent algorithms reach global solutions. Although these results are highly relevant theoretically, they have limited practical impact, as they are only asymptotic.